# SAND: One-Shot Feature Selection with Additive Noise Distortion

### v1

@@ Appendix C Benchmarking on real-world multi-spectral image dataset
The “MSI Grain” dataset was acquired using active LED light control across 15 narrow spectral bands ranging from 360 nm to 940 nm. It contains spectral data for four distinct classes: Lentils, Wood, Stones, and Background. The goal is to leverage multi-spectral imaging to classify these visually similar classes, which are difficult to distinguish with the naked eye or conventional imaging. The dataset consists of 50,000 spectra per class, extracted from several multi-spectral image cubes (see Figure 6). Each spectrum corresponds to a 15-dimensional pixel within the image, representing its spectral signature across the measured wavelengths. The dataset intentionally includes only spectral information, excluding spatial context. This design choice ensures that classification relies solely on the material’s spectral response rather than its shape. The dataset is available in the code repository at https://anonymous.4open.science/r/SAND-6BB1.
@@
Method SA LLY GL SL SAND
[…]

### v2

@@ Appendix C Benchmarking on real-world multi-spectral image dataset
The “MSI Grain” dataset was acquired using active LED light control across 15 narrow spectral bands ranging from 360 nm to 940 nm. It contains spectral data for four distinct classes: Lentils, Wood, Stones, and Background. The goal is to leverage multi-spectral imaging to classify these visually similar classes, which are difficult to distinguish with the naked eye or conventional imaging. The dataset consists of 50,000 spectra per class, extracted from several multi-spectral image cubes (see Figure 6). Each spectrum corresponds to a 15-dimensional pixel within the image, representing its spectral signature across the measured wavelengths. The dataset intentionally includes only spectral information, excluding spatial context. This design choice ensures that classification relies solely on the material’s spectral response rather than its shape. The dataset is available in the code repository at https://github.com/csem/SAND.
@@
Method SA LLY GL SL STG SAND
Accuracy 0.917 $\pm\,$ 0.003 0.917 $\pm\,$ 0.001 0.916 $\pm\,$ 0.002 0.918 $\pm\,$ 0.002 0.911 $\pm\,$ 0.005 0.919 $\pm\,$ 0.002
All Features Accuracy 0.924 $\pm\,$ 0.001